OpenStudy (anonymous):

State the value of k that makes each expression a perfect square trinomial. Then, write the trinomial as the square of a binomial: p^2 - 5p + k.

hero (hero):

p^2 - 5p + 25/4 (p - 5/2)^2

OpenStudy (anonymous):

why is it 25/4...?

OpenStudy (anonymous):

the answer is supposed to be: 6.25

OpenStudy (anonymous):

i believe they are the same number

OpenStudy (anonymous):

it is the answer but why is it 25/4?

OpenStudy (anonymous):

it is \[\frac{25}{4}\] because \[(\frac{5}{2})^2=\frac{25}{4}\]

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

and it is \[\frac{5}{2}\] because you started with \[x^2-5x+k\]

OpenStudy (anonymous):

ohh, i get it now thank you!

OpenStudy (anonymous):

Okay so here's what I did: \[p ^{2} - 5p + k =0\] coefficient of p = -5 coefficient of p / 2 = -5/2 (coefficient of p/2)^2 = (-5/2)^2 = 25.4 \[p ^{2} - 5p + 25/4 = -k + \left(\begin{matrix}25 \\ 4\end{matrix}\right)\]

OpenStudy (anonymous):

= \[(p - 5/2)^{2} = -25/4\] = \[p - 5/2 = +- (\sqrt{25}/\sqrt{4}\]

OpenStudy (anonymous):

p = -5/3 +- 5/2 Help!